OpenStudy (anonymous):

Obtain, in ascending powers of x, the first four terms in the expansion of (2-(x/2))^7. Hence find the value of (1.995)^7 correct to four decimal places. I have been able to do the first part, but am confused about the second part

OpenStudy (anonymous):

u in o levels doing add maths ???

OpenStudy (anonymous):

a binomial theorem question

OpenStudy (anonymous):

to do the second part 2-x/2=1.995 find x and put in place of x in the expansion required in the first part

OpenStudy (anonymous):

0.005 = 0.01/2

OpenStudy (anonymous):

yeah, IGCSE maths

OpenStudy (anonymous):

oh !!!

OpenStudy (anonymous):

i hope u get it

OpenStudy (anonymous):

the second part is quite simple. 1.995^7 can be rewritten in the expression given for the first part. all you have to do is, 1.995=2-x/2 find x. and then solve it using the expansion you got for the first part.

OpenStudy (anonymous):

so if my expansion is 128 - 2242x + 168x^2 - 70x^3.... then i just put 0.01 in place of x, and i should get it?

OpenStudy (anonymous):

yes u should get it

OpenStudy (anonymous):

that is correct.

OpenStudy (anonymous):

u r from .......

OpenStudy (anonymous):

india i guess

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

oh!!!! i see